# A Comprehensive Survey on Network Traffic Synthesis: From Statistical Models to Deep Learning

### v1

@@ IV-F2 K-Nearest Neighbors
Classifying network traffic can be a challenging task. Dataset imbalances, where categories are not evenly represented, can lead to biased classification results, particularly when using techniques that favor majority classes. [32]. Popular methods to balance the class distribution include random undersampling and oversampling techniques. However, these methods have notable drawbacks. Random undersampling can inadvertently discard valuable data points, while oversampling by replication may introduce bias and increase the risk of overfitting. [32, 153]. By considering the above issues, instead of over-sampling by replication, [27] introduced the Synthetic Minority Over-sampling Technique (SMOTE) to over-sample by using synthetic data. Synthetic samples are generated by operating on the feature space of minority samples and their k-nearest neighbors.
@@
In a similar effort, [153] investigates the sampling techniques used for addressing imbalanced data by classifying encrypted and unencrypted traffic. Since the unencrypted traffic is much higher than the encrypted traffic, it results in a dataset imbalance. Apart from the earlier mentioned methods, the study also investigates the use of Condensed Nearest Neighbor for undersampling and SMOTE-SVM for oversampling.
IV-G Summary
While statistical methods provided a valuable foundation for network traffic modeling and synthesis, their limitations became increasingly apparent with the discovery of self-similarity, SRD, LRD, non-stationarity, and non-linear properties in network traffic ([117, 179]). These characteristics, along with the growing complexity of modern networks, made it increasingly difficult to accurately model network traffic using traditional statistical approaches. Moreover, the abundance of available traffic data suggested that machine learning techniques, with their ability to learn complex patterns from large datasets, might be more suitable and efficient for modeling network traffic.
V Deep learning based generation
Figure 5: Trends in research papers highlighting the focus on various deep learning models over time.
This section reviews the main deep learning models used for synthetic network traffic generation: VAEs, GANs, DMs, and Transformers. Fig. 5 shows that GANs were the most widely used model in 2022, with 24 publications. Since then, there has been growing interest in Transformers and DMs, with publications on these models increasing from 2 in 2022 to 10 in 2024. This shift indicates a move towards more advanced models like DMs and Transformers (e.g., GPT), due to their ability to generate higher-quality data and better capture the time-based patterns in network traffic.
The remainder of this section provides a summary of the studies categorized by each model type. Initially, the fundamental architecture of each model is presented, followed by an analysis of how these models have been adapted for network traffic generation in various application scenarios. Furthermore, a brief discussion is included on the challenges associated with utilizing these models for network traffic data. In addition, similar to Section IV, each study is summarized in a tabular format, reporting the application scenario, model and its adaptations, data type, and artifact availability.
V-A Variational AutoEncoder (VAE)
@@
Vanilla VAE Kakkavas et al.[66] Generate network traffic matrices for network management Vanilla VAE Data link Traffic metrics (total bytes/5 min) Yes
Conditional VAE (CVAE) Aceto et al.[1] Attack/Intrusion detection CVAE Transport Packet level Yes
Liu et al.[90] Attack/Intrusion detection Vanilla VAE, CVAE Transport, Network, Data link Packet level No
VAE with Sequential Models Xiao et al.[161] Modeling QoS VAE + LSTM Transport, Network Packet level traffic, QoS data (e.g., delay/loss) No
Meslet et al.[100] Generic network traffic generation VAE + RNN + GMM Transport, Data link Packet and Flow level Yes
Miscellaneous approaches Bano et al.[16] Generic network traffic generation DNN-based noise remover + LSTM + CVAE Transport, Network, Data link Packet level No
The fundamental block of VAE is autoencoders[59, 75]. Autoencoder is a neural network architecture designed to learn efficient data representations, consisting of an encoder that compresses input data into a lower-dimensional latent space and a decoder that reconstructs the original input from this compressed representation. While effective for dimensionality reduction and feature learning, traditional autoencoders lack a probabilistic framework for generating new data. The VAE, introduced in [75] in 2013, built upon autoencoders by incorporating principles from variational inference[19], which is shown in Fig. 6. In this paper, we denote this basic VAE as vanilla VAE.
Instead of encoding inputs to arbitrary points in the latent space, the VAE’s encoder outputs parameters of a probability distribution, typically Gaussian, for each latent dimension. This probabilistic encoding allows for sampling from the latent space, enabling the generation of new, unseen data. The VAE’s decoder then learns to reconstruct the input from these sampled latent representations. This architectural shift, combined with a modified loss function that balances reconstruction quality with the regularity of the latent space (using the Kullback-Leibler divergence), transformed the deterministic autoencoder into a powerful generative model capable of both compression and novel data synthesis.
In literature, we have observed that this basic VAE is often combined with other models to enhance synthetic output. For example, to maintain long-term relationships in network time-series data, RNN or its variations (e.g., LSTM) [161, 100] are integrated with VAE. Another varient of VAE is Conditional VAE (CVAE)  [16, 1], which introduces conditional labels or attributes into the model to guide the generation process. Along these variations, VAE have been used in early studies for network traffic generation. Table IV provides a summary of existing work, with detailed discussions of each study presented next.
@@ V-A1 Vanilla VAE
V-A2 Conditional VAE (CVAE)
CVAE are used for addressing the class imbalance issue, particularly in applications such as intrusion detection systems (IDS), where the adversarial action can be less prevalent compared to legitimate users. In [1], the authors provide class labels as the conditional input when training CVAE for generating packet-level data. They propose an effective binning approach of the input samples to represent the distribution of characteristics. This binning process has helped the stable CVAE training process, while reducing the impact of noise samples through aggregation of feature values at the bin level. Also, it helps preserve the privacy of the original because the reverse mapping of bins and original data is infeasible.
@@ V-A2 Conditional VAE (CVAE)
V-A3 VAE with sequential models
To address problems such as traffic generators not being able to adapt to all types of protocols and not generating additional information like Signal to noise ratio (SNR) authors of  [100] developed an approach named Network Clustering Sequential Traffic Generation (NeCSTGen). The authors combine VAE with RNN to generate live traffic that reproduces the original behaviour at the packet, flow, and aggregate levels. Once the VAE projects the packet features into a latent space, where packets with similar features are clustered using a Gaussian Mixture Model (GMM). The RNN is then employed to ensure coherence between consecutive packets within the same flow by learning and generating transitions between these clusters, maintaining short-term dynamics and flow-level consistency.
In [161], the authors developed LSTM-enhanced VAE to model the Quality of Service (QoS). The QoS inference problem is defined as inferring the exact value of QoS metrics at each moment given real-time traffic loads. The LSTM module is used to extract fine-grained traffic features as a condition to the VAE and the VAE module learns to reconstruct the QoS distribution from the traffic features.
V-A4 Miscellaneous approaches
In  [16], a CVAE combined with LSTM is utilized for generating synthetic network traffic and imputing missing data within traffic streams. The data generation process involves two stages. First, a DNN model is applied as a preprocessing step to remove noisy packets and retain only legitimate traffic. Second, the legitimate traffic is used to train the CVAE, with conditional inputs (e.g., traffic attributes or labels) provided at the decoder. LSTM models are employed in both the encoder and decoder to capture temporal relationships within the traffic data. The generated data aligns with the statistical distribution of the original traffic, preserving real-world dynamics even in the presence of missing input data.
V-A5 Limitations in VAE based traffic generation
@@ V-A6 Summary
Vanilla GAN Nukavarapu et al. [109] Generic network traffic generation Vanilla GAN Transport, Network Packet level No
Cheng et al. [30] Generic network traffic generation Vanilla GAN Application Bytes in hexadecimal notation No
@@
Noel et al. [107] Generic network traffic generation for darknet data Vanilla GAN Transport, Network Packet level No
@@
Xiao et al. [162] Generic network traffic generation for unbalanced distributed data Vanilla GAN, Federated learning Transport, Network Packet level No
Guo et al. [55] Improve ML accuracy/class imabalance Vanilla GAN Application Bytes in hexadecimal notation No
@@
Shahriar et al. [133] Handle class imbalance, Attack/Intrusion detection Vanilla GAN Transport, Network Packet level No
@@
WGAN Ring et al. [128] Generic network traffic generation WGAN Transport Flow level No
Fathi et al. [48] Traffic obfuscation WGAN Transport, Network Flow level No
Zhang et al. [175] Traffic obfuscation WGAN Transport Packet level No
Mozo et al. [103] Attack/Intrusion detection WGAN Transport, Network Flow level No
@@
Kattadige et al. [68] Improve ML accuracy/class imbalance WGAN Transport Packet level Yes
@@
Madarasingha et al. [94] Improve ML accuracy/class imbalance WGAN Transport Packet level Yes
@@
Sivaroopan et al.[137] Improve ML accuracy/class imbalance WGAN Transport, Network Packet level Yes
Roy et al.[130] Generic RF signal generation WGAN Physical Raw signal No
@@
CGAN Jiang et al. [63] Generic network traffic generation CGAN Transport, Network Flow level No
Wang et al. [156] Improve ML accuracy/class imbalance CGAN Application Bytes in hexadecimal notation No
@@
Kotal et al. [78] Improve ML accuracy/class imbalance CGAN Transport, Network Flow level No
@@
Almasre et al. [5] IoT traffic generation for cybersecurity applications CGAN Transport, Network Packet and Flow level Yes
@@
Sequential Model Variants Lin et al. [87] Handle class imbalance / Privacy-preserved data sharing GAN, LSTM Transport Flow level Yes
@@
Yin et al. [171] Handle class imbalance / Privacy-preserved data sharing GAN, LSTM Transport, Network Packet and Flow level Yes
@@
Zhang et al. [173] Improve ML accuracy/class imbalance GAN, Transformer Transport, Network Flow level No
@@
Shahid et al. [132] Attack/Intrusion detection Autoencoder, GAN, LSTM Transport Packet level No
Qu et al.[123] Attack/Intrusion detection GAN, LSTM Transport Packet level No
@@
Misc. Approaches Anande et al. [8] Generic network traffic generation Vanilla GAN, CTGAN, Copula GAN Network, Transport Flow level No
@@
Meddahi et al. [97] Generic network traffic generation for SIP protocol DCGAN Application Bytes in hexadecimal notation Yes
@@
Yang et al. [168] Generate network traffic Swing GAN, Autoencoders Transport, Network Flow level No
@@
Kim et al. [73] Simulation tool development GAN-based tool Transport, Network, Data-link Packet level No
@@
Dowoo et al.[45] Handle data shortage in cyber-domain Graph GAN, Image GAN, Sequence GAN, Makrov models Transport, Network Packet level No
Khan et al.[70] Handle class imbalance BEGAN N/A N/A No
V-B Generative Adversarial Networks (GAN)
@@
GANs, introduced by Goodfellow et al. [52] in 2014, represented a groundbreaking approach in generative modeling. As shown in Fig. 7, GANs consist of two competing neural networks: a generator and a discriminator. The generator’s objective is to capture the data distribution and generate realistic samples, while the discriminator’s role is to differentiate between real samples from the training data and fake samples produced by the generator. The two models are trained simultaneously in anadversarial process. The generator attempts to produce data that closely resembles the real data, aiming to fool the discriminator. Meanwhile, the discriminator works to improve its accuracy in distinguishing between genuine data and the data generated by the generator. This adversarial setup results in a continuous improvement loop: as the generator becomes better at creating realistic data, the discriminator must also become more adept at distinguishing between real and fake data. The training procedure can be described as a min-max two-player game, where the generator seeks to maximize the probability the discriminator makes a mistake, while the discriminator attempts to minimize the mistakes.
Although GANs were initially used to generate images, researchers began applying them to time-series data. This shift has enabled the generation of network traffic, as the structure and characteristics of network traffic closely align with those of time-series data. We identify five main GAN-based model architectures in network traffic generation literature.
@@
V-B1 Vanilla GAN
The basic GAN model, which comprises of only generator and discriminator modules, has shown promising results in maintaining network traffic properties, despite the sequential nature they present, in network traffic generation. The authors in [109] focused on creating a protocol-agnostic framework for packet generation developing a platform named MiragePkt. In their approach, network packets were modeled as sequences of bytes. To create these sequences they leverage packet tokenization techniques, and GANs were employed to understand the conditional distribution of bytes within these sequences. To properly understand the properties of sequence, MiragePkt consists of 1D CNN layers. They further apply post-processing techniques to convert the output byte data from the GAN model to a hex stream.

### v2

@@ III-F2 K-Nearest Neighbors (KNN)
Classifying network traffic can be a challenging task. Dataset imbalances, where categories are not evenly represented, can lead to biased classification results, particularly when using techniques that favor majority classes. [36]. Popular methods to balance the class distribution include random undersampling and oversampling techniques. However, these methods have notable drawbacks. Random undersampling can inadvertently discard valuable data points, while oversampling by replication may introduce bias and increase the risk of overfitting. [36, 170]. By considering the above issues, instead of over-sampling by replication, [31] introduced the Synthetic Minority Over-sampling Technique (SMOTE) to over-sample by using synthetic data. Synthetic samples are generated by operating on the feature space of minority samples and their KNN.
@@
In a similar effort, [170] investigates the sampling techniques used for addressing imbalanced data by classifying encrypted and unencrypted traffic. Since the unencrypted traffic is much higher than the encrypted traffic, it results in a dataset imbalance. Apart from the earlier mentioned methods, the study also investigates the use of Condensed Nearest Neighbor for undersampling and SMOTE with a Support Vector Machine (SMOTE-SVM) for oversampling.
Limitations: The studies reviewed indicate that KNN and SMOTE-based approaches are inherently dependent on feature-space structure and data quality. As observed in [36, 170], inappropriate sampling strategies can introduce bias or overfitting, and these methods do not model temporal traffic dynamics. Consequently, their applicability is limited to data balancing rather than traffic behavior synthesis.
Summary: In summary, KNN and SMOTE-based methods address a complementary problem in network traffic analysis: mitigating class imbalance. The reviewed works demonstrate their effectiveness in improving classification performance for intrusion detection and encrypted traffic analysis [36, 170]. These methods function as augmentation techniques that enhance downstream learning tasks rather than as generative traffic models.
III-G Summary
While statistical methods provided a valuable foundation for network traffic modeling and synthesis, their limitations became increasingly apparent with the discovery of self-similarity, SRD, LRD, non-stationarity, and non-linear properties in network traffic ([136, 206]). These characteristics, along with the growing complexity of modern networks, made it increasingly difficult to accurately model network traffic using traditional statistical approaches. Moreover, the abundance of available traffic data suggested that ML techniques, with their ability to learn complex patterns from large datasets, might be more efficient for modeling network traffic.
IV DL based generation
Figure 6: Trends in research papers highlighting the focus on various DL models over time.
This section reviews the main DL models used for synthetic network traffic generation: VAEs, GANs, DMs, and Transformers. Fig. 6 shows that GANs were the most widely used model in 2022, with 24 publications. Since then, there has been growing interest in Transformers and DMs, with publications on these models increasing from 2 in 2022 to 10 in 2024. This shift indicates a move towards more advanced models like DMs and Transformers (e.g., GPT), due to their ability to generate higher-quality data and better capture the time-based patterns in network traffic.
The remainder of this section provides a summary of the studies categorized by each model type. Initially, the fundamental architecture of each model is presented, followed by an analysis of how these models have been adapted for network traffic generation in various application scenarios. Furthermore, a brief discussion is included on the challenges associated with utilizing these models for network traffic data. In addition, similar to Section III, each study is summarized in a tabular format, reporting the application scenario, model and its adaptations, data type, and artifact availability.
IV-A Variational Autoencoder (VAE)
@@
Vanilla VAE [78] Generate network traffic matrices for network management Vanilla VAE Data link Traffic metrics (total bytes/5 min) Yes
[66] Food sensing Vanilla VAE Physical RF signals Yes
[107] Estimating downlink wireless channel for MIMO Vanilla VAE Physical RF signals No
[193] WI-FI fall detection Vanilla VAE Physical WI-FI No
[161] Indoor localization fingerprinting Vanilla VAE Physical RSSI No
Conditional VAE (CVAE) [1] Attack/Intrusion detection CVAE Transport Packet level Yes
[105] Attack/Intrusion detection Vanilla VAE, CVAE Transport, Network, Data link Packet level No
[18] Channel estimation CVAE Physical CSI data No
VAE with Sequential Models [183] Modeling QoS VAE + LSTM Transport, Network Packet level, QoS data (e.g., delay/loss) No
[116] Generic network traffic generation VAE + RNN + GMM Transport, Data link Packet and Flow level Yes
Misc. approaches [17] Generic network traffic generation DNN-based noise remover + LSTM + CVAE Transport, Network, Data link Packet level No
Autoencoder is a neural network architecture designed to learn efficient data representations, consisting of an encoder that compresses input data into a lower-dimensional latent space and a decoder that reconstructs the original input from this compressed representation. While effective for dimensionality reduction and feature learning, traditional autoencoders lack a probabilistic framework for generating new data. The VAE, introduced in [89] in 2013, built upon autoencoders by incorporating principles from variational inference[22], which is shown in Fig. 7. We denote this basic VAE as vanilla VAE.
Instead of encoding inputs to arbitrary points in the latent space, the VAE’s encoder outputs parameters of a probability distribution, typically Gaussian, for each latent dimension. This probabilistic encoding allows for sampling from the latent space, enabling the generation of new, unseen data. The VAE’s decoder then learns to reconstruct the input from these sampled latent representations. This architectural shift, combined with a modified loss function that balances reconstruction quality with the regularity of the latent space, transformed the deterministic autoencoder into a powerful generative model capable of both compression and novel data synthesis. In literature, we have observed that this basic VAE is often combined with other models to enhance synthetic output. For example, to maintain long-term relationships in network time-series data, Recurrent Neural Network (RNN) or its variations (e.g., Long short-term memory [LSTM]) [183, 116] are integrated with VAE. Another variant of VAE is Conditional VAE (CVAE) [17, 1], which introduces conditional labels or attributes into the model to guide the generation process. Table IV provides a summary of existing work.
@@ IV-A1 Vanilla VAE
RF-EATS [66], a noninvasive system leveraging VAE, uses passive backscatter tags and near-field coupling to sense food and liquids inside closed containers without contact or calibration. FIRE [107], leveraging VAE, proposes an end-to-end machine learning system for accurate downlink channel estimation in massive MIMO without requiring feedback from client devices. FIRE learns a latent representation from uplink channels to generate downlink estimates, offering an interpretable, efficient, and scalable solution for real-world deployments. In [193], the authors present FallDar, a Wi-Fi-based DL fall detection system that addresses key challenges of environmental, motion, and user diversity. FallDar uses body speed as a robust feature, generates diverse fall scenarios with VAE, and incorporates a user identification network to enable accurate detection without needing user-specific fall data. In [161], the authors address the high cost and effort of constructing offline databases for fingerprinting-based indoor localization by leveraging VAE. Using real-world RSSI data, the VAE learns the distribution of RSSI to generate synthetic values, mitigating the challenges posed by its inherent instability. This approach aims to enhance the efficiency and reliability of DL-based indoor localization systems.
IV-A2 Conditional VAE (CVAE)
CVAE are used for addressing the class imbalance issue, particularly in applications such as IDS, where the adversarial action can be less prevalent compared to legitimate users. In [1], the authors provide class labels as the conditional input when training CVAE for generating packet-level data. They propose an effective binning approach of the input samples to represent the distribution of characteristics. This binning process has helped the stable CVAE training process, while reducing the impact of noise samples through aggregation of feature values at the bin level. Also, it helps preserve the privacy of the original because the reverse mapping of bins and original data is infeasible.
@@ IV-A2 Conditional VAE (CVAE)
In [18], a CVAE is proposed for data-driven channel estimation by modeling the unknown channel distribution as a conditional Gaussian parameterized by conditional moments. The authors introduce three VAE-based estimators, showing that they can approximate the minimum mean square error (MMSE) estimator, with practical variants requiring full channel knowledge only during training or not at all.
IV-A3 VAE with sequential models
To address problems such as traffic generators not being able to adapt to all types of protocols and not generating additional information like Signal to noise ratio (SNR) authors of [116] developed an approach named Network Clustering Sequential Traffic Generation (NeCSTGen). The authors combine VAE with RNN to generate live traffic that reproduces the original behaviour at the packet, flow, and aggregate levels. Once the VAE projects the packet features into a latent space, where packets with similar features are clustered using a Gaussian Mixture Model (GMM). The RNN is then employed to ensure coherence between consecutive packets within the same flow by learning and generating transitions between these clusters, maintaining short-term dynamics and flow-level consistency. In [183], the authors developed LSTM-enhanced VAE to model the QoS. The QoS inference problem is defined as inferring the exact value of QoS metrics at each moment given real-time traffic loads. The LSTM module is used to extract fine-grained traffic features as a condition to the VAE and the VAE module learns to reconstruct the QoS distribution from the traffic features.
IV-A4 Miscellaneous approaches
In [17], a CVAE combined with LSTM is utilized for generating synthetic network traffic and imputing missing data within traffic streams. The data generation process involves two stages. First, a DNN model is applied as a preprocessing step to remove noisy packets and retain only legitimate traffic. Second, the legitimate traffic is used to train the CVAE, with conditional inputs (e.g., traffic attributes or labels) provided at the decoder. LSTM models are employed in both the encoder and decoder to capture temporal relationships within the traffic data. The generated data aligns with the statistical distribution of the original traffic, preserving real-world dynamics even with missing data.
IV-A5 Limitations in VAE based traffic generation
@@ IV-A6 Summary
Vanilla GAN [125] Generic network traffic generation Vanilla GAN Transport, Network Packet level No
[33] Generic network traffic generation Vanilla GAN Application Payload No
@@
[123] Generic network traffic generation for darknet data Vanilla GAN Transport, Network Packet level No
@@
[184] Generic network traffic generation Vanilla GAN, FL Transport, Network Packet level No
[65] Improve ML accuracy/class imbalance Vanilla GAN Application Payload No
@@
[153] Handle class imbalance, Attack/Intrusion detection Vanilla GAN Transport, Network Packet level No
@@
WGAN [147] Generic network traffic generation WGAN Transport Flow level No
[55] Traffic obfuscation WGAN Transport, Network Flow level No
[202] Traffic obfuscation WGAN Transport Packet level No
[119] Attack/Intrusion detection WGAN Transport Flow level No
@@
[81] Improve ML accuracy/class imbalance WGAN Transport Packet level Yes
@@
[110] Improve ML accuracy/class imbalance WGAN Transport Packet level Yes
@@
[157] Improve ML accuracy/class imbalance WGAN Transport, Network Packet level Yes
[150] Generic RF signal generation WGAN Physical Raw signal No
[16] Wideband channel estimation WGAN Physical mmWave No
@@
CGAN [75] Generic network traffic generation CGAN Transport, Network Flow level No
[176] Improve ML accuracy/class imbalance CGAN Application Payload No
@@
[92] Improve ML accuracy/class imbalance CGAN Transport, Network Flow level No
@@
[5] IoT traffic generation for cybersecurity applications CGAN Transport, Network Packet & Flow Yes
[54] Human activity recognition ACGAN Physical RF signal No
[50] mmWave MIMO channel estimation CGAN with FL Physical MIMO matrix No
[187] Anomaly detection CGAN Physical Wireless No
@@
Seq. Model Variants [102] Class imbalance / Privacy-preserved data sharing GAN, LSTM Transport Flow level Yes
@@
[197] Class imbalance / privacy-preserved data sharing GAN, LSTM Transport, Network Packet & Flow Yes
@@
[200] Improve ML accuracy/class imbalance GAN, Transformer Transport, Network Flow level No
@@
[152] Attack/Intrusion detection Autoencoder, GAN, LSTM Transport Packet level No
[142] Attack/Intrusion detection GAN, LSTM Transport Packet level No
@@
Misc. Approach [8] Generic network traffic generation CTGAN, Copula GAN Network, Transport Flow level No
@@
[113] Generic network traffic generation for SIP protocol DCGAN Application Payload Yes
@@
[191] Generate network traffic Swing GAN, Autoencoders Transport, Network Flow level No
@@
[87] Simulation tool development GAN-based tool Transport, Network, Data-link Packet level No
@@
[51] Handle data shortage in cyber-domain Graph/Image/Sequence GAN, Makrov models Transport, Network Packet level No
[84] Handle class imbalance BEGAN N/A N/A No
IV-B Generative Adversarial Networks (GAN)
@@
GANs, introduced by Goodfellow et al. [61] in 2014, represented a groundbreaking approach in generative modelling. As shown in Fig. 8, GANs consist of two competing neural networks: a generator and a discriminator. The generator’s objective is to capture the data distribution and generate realistic samples, while the discriminator’s role is to differentiate between real samples from the training data and fake samples produced by the generator. The two models are trained simultaneously in an adversarial process. The generator attempts to produce data that closely resembles the real data, aiming to fool the discriminator. Meanwhile, the discriminator works to improve its accuracy in distinguishing between genuine data and the data generated by the generator. This adversarial setup results in a continuous improvement loop: as the generator becomes better at creating realistic data, the discriminator must also become more adept at distinguishing between real and fake data. The training procedure can be described as a min-max two-player game, where the generator seeks to maximize the probability the discriminator makes a mistake, while the discriminator attempts to minimize the mistakes. Although GANs were initially used for image data, researchers began applying them to time-series data including network traffic data. We identify five main GAN-based model architectures in this domain.
@@
IV-B1 Vanilla GAN
The basic GAN model, which comprises of only generator and discriminator modules, has shown promising results in maintaining network traffic properties, despite the sequential nature they present, in network traffic generation. The authors in [125] focused on creating a protocol-agnostic framework for packet generation developing a platform named MiragePkt. In their approach, network packets were modeled as sequences of bytes. To create these sequences they leverage packet tokenization techniques, and GANs were employed to understand the conditional distribution of bytes within these sequences. To properly understand the properties of sequence, MiragePkt consists of 1D Convolutional Neural Network (CNN) layers. They further apply post-processing techniques to convert the output byte data from the GAN model to a hex stream.